write an equation for the line 3x+6y=9 in the form y=mx+b
Answer by stanbon(75887)

write an equation for the line 3x+6y=9 in the form y=mx+b
----------
3x+6y = 9
Divide thru by 3
---------------------------
x + y = 3
Solve for "y":
---------------------------
y = -x+3
-------------
Cheers,
Stan H.